# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### 14 Oct 2024 v1

@@ 6.3 Hardware
In response to the steep hardware requirement, many creators rely on cloud GPU services. Based on our analysis of the links shared on MrDeepFakes, the most commonly discussed is Google Colab (a hosted version of Jupyter Notebook that provides users access to GPUs via a freemium mode), accounting for 92 links, followed by Paperspace (26) and AWS (10). As one community member in 2019 noted: “Colab works so much better than my own PC. Do not bother trying to create fakes on a Radeon RX 580, what took me a day is done in an hour using Colab.” However, as the community discusses, cloud GPU providers pose unique challenges, including the nondeterminism of provisioned GPUs, unexpected runtime disconnections, and threats to users’ privacy (e.g., not accepting cryptocurrency payments, server-side scanning).
Furthermore, cloud GPU users are subject to provider ToS; for instance, in May 2022, Google banned the use of Colab for training deepfakes [63], which immediately caused ripples in the community:
@@
Aligned with prior explorations of the efficacy of deplatforming abusive actors [7, 30, 23, 25] and the pursuit of target hardening [50], we anecdotally find that while major platform policy changes do not exhaustively eradicate abuse, they do significantly increase the cost of deepfake creation; such changes in the deepfake value chain may be enough to turn away less capable (e.g., financially, technically) creators and stem the growth of this ecosystem.
@@ 6.3.1 Google Colab Deepfake Ban
In May 2022, Google banned using Colab notebooks for training models to generate deepfakes [63]. In the following month, MrDeepFakes members began to respond with reports of service interruptions, calls for help to the community, and speculations of how the ban was implemented and enforced:
@@
Members noted that Colab was flagging deepfake-related notebooks by detecting use of the DFL library. Almost immediately, the MrDeepFakes community sought to identify opportunities to circumvent the ban. Several members shared trivial yet effective solutions that obfuscated the source of deepfake code by forking the original DFL repository and then cloning the fork, reducing traceability to the banned DFL repository:
@@
While discussion of Colab has slowed since the ban, continued reports of service interruptions and iterative development of ban evasion solutions illustrate the cat-and-mouse game between Google and the community. The community concluded that Google’s detection of DFL naively relies on a blocklist, resulting in recommendations to “not share [their DFL] Github fork and try to give [it] a unique name.” The risk of informing Google’s deepfake detection has also led to some members seeking to move ban evasion support to private channels, asking others to “please DM if [they] know of any tricks.” As of July 2024, it is unclear whether community solutions still work today; some users report that they are able to use Colab, whereas others report errors.
@@ 6.4 Community Support
The MrDeepFakes community appears to be responsive and welcoming to newcomers. 80.6% of question threads on MrDeepFakes received a reply, with the median time to receive a reply of 3.4 hours ($\mu=5.8$ days, $\sigma=39.3$ days), which is similar to Stack Overflow, where 80% of posts receive a response and posts have a median response time of 40 minutes ($\mu=11.7$ days, $\sigma=99.2$ days). While MrDeepFakes’s median time to response is longer than that of Stack Overflow, we suspect that this is largely a function of community size; as of August 2024, the Stack Overflow community had 26M members [13], as opposed to the 611K members of MrDeepFakes. On the other hand, the mean response time on MrDeepFakes is half that of Stack Overflow, illustrating that the distribution of response times on MrDeepFakes is much tighter. We also find that 72% of questions on MrDeepFakes receive a response in less than three days. Therefore, while members of MrDeepFakes may not receive instantaneous responses, they can reliably expect to receive technical support within a reasonable time frame. Newcomers with accounts under 45 days old start 39.3% of all question threads
Unlike Stack Overflow, which is known to enforce a “no pleases or thank yous” rule [13], the MrDeepFakes community does not have any rules regarding the tone or delivery of questions in these forum sections. Anecdotally, we observe that the community encourages pleasantries and more personal connections between members, in part, perhaps because enterprising members are looking to attract clientele:
PM me if you’ll need more help, I offer personal courses…Or read my guide, it covers all the basics if you never used DFL to make fakes.
@@ 7 Discussion and Conclusion
Grounded in our data, we present the following contributions: (1) we measure the growth of MrDeepFakes and the targets of deepfake videos; (2) we characterize the participants of the deepfake market and analyze the prices of bespoke sexual deepfakes; and (3) we document the required components of deepfake creation and how the community supports members in the process. We synthesize the implications of our findings on disincentivizing deepfake NCII production and mitigating its harms below.
@@ 7.1 MrDeepFakes as a Community
MrDeepFakes has grown immensely, drawing over 1.5B views of 43K sexual deepfake videos. The scale of site engagement highlights that consumption and discussion of sexual deepfakes occurs openly online and has become mainstream. Its growth has led to the maturation of its profit-driven marketplace for sexual deepfakes. As the market continues to develop, we suspect we will observe even more seller specialization than currently exists, similar to other specialized labor markets [24]—for example, in response to tightening margins on celebrity deepfakes, one MrDeepFakes member suggested that you could make money by applying existing tools to do “many things” beyond celebrity fakes (Section 5.3).
Deepfakes as Art.
We found that the MrDeepFakes community is welcoming and responsive; many of its members and leadership share a value system that positions deepfake creators as artists and the sexual deepfakes they produce as artistic contributions that benefit their community. This novel incentive differentiates deepfake creators as a class of actors separate from previously studied cybercrime and online abuse attackers [56, 10]; thus, we identify two immediate areas for future work: (1) understanding to what degree deepfake creators perceive the harms (and the “benefits”) of their work and (2) investigating potential interventions for deepfake creators, as prior work has found effective in combating misinformation [21], antisocial behaviors [8], and poor cybersecurity behaviors [65].
@@
Core members are technically skilled and willing to share their technical expertise on deepfake creation. Contrary to the beliefs of open source developers of a SFW deepfake tool studied by Widder et al. [62], denying technical support does not appear a viable mitigation for the abuse of deepfake creation tools. Furthermore, significant sharing of academic work within the community shows that its members closely follow and utilize academic research, highlighting the dual-use potential for many computing innovations. Similar to the OSS community, many instances of academic research have used technology neutrality and inevitability to justify the publication and dissemination of innovative developments. We acknowledge that the solution does not lie in silencing research findings or technological innovation, but we argue that researchers and developers should consider principles of “target hardening” in their own work to (1) carefully choose how to distribute their work and when possible, (2) preemptively design safeguards to promote the intended use and discourage the abuse of their research.
@@ 7.2 Deepfake Abuse Mitigation is Adversarial
Technical solutions to preventing deepfake creation are limited, as evidenced by adversarial efforts of MrDeepFakes members to circumvent existing efforts by cloud GPU providers. While identifying platform-centered choke points can reduce harm, as the security community has concluded for numerous abuse ecosystems [14, 17, 26], the endless emergence of alternative “bulletproof” service providers has limited the efficacy of policy-centered mitigation (i.e., deplatforming). In the face of complex security problems with no tractible, perfect solutions, we acknowledge the impact and importance of target hardening [50]; however, we identify a promising approach to more effectively stem deepfake abuse through legislation.
@@
A MrDeepFakes staff member shared that enactment of legislation internationally has already begun to make monetizing sexual deepfakes more difficult, implying that laws have significant sway over the deepfake economy (Section 5.3). In February 2024, following a push for federal legislation protecting against deepfake abuse [53], forum members asked staff for their reflections on the implications of federally outlawing deepfake abuse, to which a staff member replied:
@@
This suggests that laws are powerful enough to disrupt behaviors of even highly skilled, core members of the community. We argue that augmenting our existing mitigation approach by incorporating legislative protections has significant potential to more effectively stem deepfake abuse.
Ethics
@@
In this paper, we analyze MrDeepFakes, a publicly available platform for hosting, buying, and selling sexual deepfakes. Our work is in the spirit of significant prior work by the security community [38, 58, 60, 55, 15]. However, our context is significantly different from prior work, and we carefully evaluated the ethical implications of our research. We considered many stakeholders when evaluating the ethics of our work: the MrDeepFakes platform itself, its users, depicted and discussed targets, and the researchers who studied the platform given its potentially harmful content. We discuss the ethical implications of our research choices for each stakeholder below:
The MrDeepFakes Platform.
@@
MrDeepFakes’ terms of service prohibit using site content for commercial endeavors, redistributing content, and sending users unsolicited emails, but they do not prohibit scraping or automated data collection. The site’s robots.txt only disallows crawling by the Internet Archive (User-agent: ia_archiver). All of the information we scraped from the website is accessible to anyone with an account.  [^4] Account creation is free and does not require e-mail verification; site operators have also stated that users are not obligated to provide “real” information. We acknowledge that in scraping the website, we may have increased server load and bandwidth on the part of the platform; however, we minimized the number of requests made to the platform, and our request volume is relatively low compared to the organic activity of tens of thousands of accounts that the platform services. We stress that we did not collect or store any video or image data, only collecting relevant metadata required to answer our research questions.
@@
We did consider whether to publish the name of the platform itself in this paper, as similar prior work opted to not reveal the name of the platforms in question as to not draw potential attention to its existence [58]. However, we decided to name the MrDeepFakes platform as it has been named directly in prior work [57, 47] and is the top search result for celebrity sexual deepfakes. We hope that naming this platform can enable future research and potential regulation on the issue of sexual deepfake creation.
Users of the MrDeepFakes platform.
@@
In examining user data, we collected only publicly available data, and the only potentially personally identifying information we collected was the account username and the user ID. We never attempted to deanonymize any user in our dataset and we did not interact with any community members in any fashion (e.g., via direct messages or public posts). In our report, we take careful note to not include any users’ account details when reporting quotes and opt to only report aggregate statistics to protect user privacy.
@@
To avoid drawing undue attention to any targets mentioned in the forum, we redact and paraphase any personally identifying information in the quotes we reported and elect not to name any target in the text.
@@
Given that this paper required qualitative analysis of text-based forum data, the researchers involved in these tasks were directly exposed to potentially disturbing content that might cause significant psychological or emotional harms. To mitigate these issues, the researchers had regular check-ins regarding their psychological well being, had access to therapy and counseling services, and took breaks as needed.
@@
We are committed to sharing our data and code with other researchers at academic, non-commercial institutions seeking to conduct future work. However, given the sensitive nature of the data and disturbing nature of the content, we will not make this data available publicly. We note that the site content is available on the open Internet and that motivated actors can easily access the content for themselves. However, we do not want to enable malicious actors seeking to use MrDeepFakes data to potentially harm others. We are committed to sharing our data and our codebooks with the Artifact Evaluation committee to ensure our artifacts meet the USENIX Open Science standards.
Appendix A Forum map

### 02 Jan 2025 v2

@@ 6.4 Hardware
I wish I could buy my own GPU. It’s too expensive to import to my country, but maybe one day :(
@@
the electric cost of running gpus 24/7 will be too high
@@
Cloud GPU Providers.In response to the steep hardware requirement, many creators rely on cloud GPU services. Based on our analysis of the links shared on MrDeepFakes, the most commonly discussed is Google Colab (a hosted version of Jupyter Notebook that provides users access to GPUs via a freemium mode), accounting for 92 links, followed by Paperspace (26) and AWS (10). As one community member in 2019 noted: “Colab works so much better than my own PC. Do not bother trying to create fakes on a Radeon RX 580, what took me a day is done in an hour using Colab.” However, as the community discusses, cloud GPU providers pose unique challenges, including the nondeterminism of provisioned GPUs, unexpected runtime disconnections, and threats to users’ privacy (e.g., not accepting cryptocurrency payments, server-side scanning).
Furthermore, cloud GPU users are subject to provider ToS; for instance, in May 2022, Google banned the use of Colab for training deepfakes [google-colab-ban], which immediately caused ripples in the community:
@@
We can’t make free deepfakes on Colab because Google banned it. If you have a free account your runtime will be repeatedly disconnected. Eventually they might ban you from Colab.
@@
I actually never used Runpod, so maybe check out YouTube for tutorials. Or you can try Paperspace, but it might cost more. If electricity costs keep rising then I might have to use cloud service bullshit too.
Aligned with prior explorations of the efficacy of deplatforming abusive actors [chandrasekharan2017you, levchenko2011trajectories, kanich2008spamalytics, bellingcat2023any] and the pursuit of target hardening [redmiles2024friction], we anecdotally find that while major platform policy changes do not exhaustively eradicate abuse, they do significantly increase the cost of deepfake creation; such changes in the deepfake value chain may be enough to turn away less capable (e.g., financially, technically) creators and stem the growth of this ecosystem.
@@ 6.4.1 Google Colab Deepfake Ban
In May 2022, Google banned using Colab notebooks for training models to generate deepfakes [google-colab-ban]. In the following month, users MrDeepFakes members began to respond with reports of service interruptions, calls for help to the community, and speculations of how the ban was implemented and enforced:
@@
Maybe Google is detecting DFL usage then terminating those Colab notebooks. I hope it’s temporary I don’t know any other free options.
Members noted that Colab was flagging deepfake-related notebooks by detecting use of the DFL library. Almost immediately, the MrDeepFakes community sought to identify opportunities to circumvent the ban. Several members shared trivial yet effective solutions that obfuscated the source of deepfake code by forking the original DeepFaceLab DFL repository and then cloning the fork, reducing traceability to the banned DeepFaceLab repository:
@@
DFL repository:
You don’t need to know how to code, just copy and paste the notebook cell changes. This simple fix should take you at most 30 minutes and lets you circumvent Google’s detection of DFL.
While discussion of Colab has slowed since the ban, continued reports of service interruptions and iterative development of ban evasion solutions illustrate the cat-and-mouse game between Google and the community. The community concluded that Google’s detection of DeepFaceLab DFL naively relies on a blocklist, resulting in recommendations to “not share [their DFL] Github fork and try to give [it] a unique name.” The risk of informing Google’s deepfake detection has also led to some members seeking to move ban evasion support to private channels, asking others to “please DM if [they] know of any tricks.” As of July 2024, it is unclear whether community solutions still work today; some users report that they are able to use Colab, whereas others report errors.
@@ 6.5 Community Support
The MrDeepFakes community appears to be responsive and welcoming to newcomers. 80.6% of question threads on MrDeepFakes received a reply, with the median time to receive a reply of 3.4 hours ($\mu=5.8$ days, $\sigma=39.3$ days), which is similar to Stack Overflow, where 80% of posts receive a response and posts have a median response time of 40 minutes ($\mu=11.7$ days, $\sigma=99.2$ days). While MrDeepFakes’s median time to response is longer than that of Stack Overflow, we suspect that this is largely a function of community size; as of August 2024, the Stack Overflow community had 26M members [stackexchange2024sites], as opposed to the 611K members of MrDeepFakes. On the other hand, the mean response time on MrDeepFakes is half that of Stack Overflow, illustrating that the distribution of response times on MrDeepFakes is much tighter. We also find that 72% of questions on MrDeepFakes receive a response in less than three days. Therefore, while members of MrDeepFakes may not receive instantaneous responses, they can reliably expect to receive technical support within a reasonable time frame. Newcomers with accounts under 45 days old start 39.3% of all question threads
Unlike Stack Overflow, which is known to enforce a “no pleases or thank yous” rule [stackexchange2024sites], the MrDeepFakes community does not have any rules regarding the tone or delivery of questions in these forum sections. Anecdotally, we observe that the community encourages pleasantries and more personal connections between members, in part, perhaps because enterprising members are looking to attract clientele:PM me if you’ll need more help, I offer personal courses…Or read my guide, it covers all the basics if you never used DFL to make fakes.
DM me for personal help, I offer private tutoring. Or you can read the beginners’ guides I wrote for using DFL for deepfakes.
@@
We are the only legit deepfake site with an actual community, the last thing we want is to scam our members.
@@ 7 Discussionand Conclusion
Grounded in our data, we present the following contributionsDeepfake IBSA causes and amplifies a wide range of real harms. It is commonly used to financially sextort targets, generate “revenge pornography,” and explore sexual fantasies without obtaining consent [wei2024understanding, burkell2019nothing, kshetri2023economics, okolie2023ibsa]. Development of generative AI technology has made deepfake capabilities more accessible, promoting widespread use—including by children [maiberg22clicks]. In this work (Sections 46), we: (1) we measure measured the growth of MrDeepFakes and the targets of deepfake videos; (2) we characterize characterized the participants of the deepfake market and analyze analyzed the prices of bespoke sexual deepfakes; and (3) we document documented the required components of deepfake creation and how the community supports members in the process. We synthesize the implications of our findings on disincentivizing deepfake NCII production and mitigating its harms below.
@@ 7.1 MrDeepFakes as a Community
MrDeepFakes has grown immensely over the past few years, drawing over 1.5 billion views of 43K sexual deepfake videos . as of November 2023. The scale of site engagement highlights that consumption and discussion of sexual deepfakes occurs openly online and has become mainstream . Its and occurs openly online. This growth has led to the maturation of its profit-driven marketplace for sexual deepfakes. As the market matures, we suspect we will observe even more seller specialization than currently exists, continues to develop, seller specialization will likely also further develop similar to other specialized labor markets [sunwoong1989labor]—for . The market will likely also expand. For example, in response to tightening margins on celebrity deepfakes, one MrDeepFakes member suggested that you could make money by applying existing tools to do “many things” beyond celebrity fakes (Section 5.4).
Deepfakes as Art
We also found that the
Deepfakes as Art. The MrDeepFakes community is welcoming and responsive; many of its members and leadership leaders share a value system that positions deepfake creators as artists and the sexual deepfakes they produce as artistic contributions that benefit their the community. This novel incentive unique motivation differentiates deepfake creators as a class of actors separate from previously studied cybercrime and online abuse attackers [thomas2021hatesok, chng2022hacker]; thusactors [thomas2021hatesok, chng2022hacker]. To better understand and disincentivize this class of perpetrator, we identify two immediate three areas for future work: (1) understanding to what degree deepfake creatorsperceive the harms (and the “benefits ”) first, characterizing deepfake creators’ perceptions of the benefits and harms of their workand (2) investigating potential ; second, investigating the applicability of interventions for deepfake creators, as such as targeting inaccurate meta-cognitive (i.e., second-order) beliefs and nudging, which prior work has found effective in for combating misinformation [Jahanbakhsh2024signaling], antisocial behaviors [cheng2017anyone], [jeffrey2021understanding] and poor cybersecurity behaviors [zimmerman2021nudge]. practices [zimmerman2021nudge], respectively; third, drawing on Insoll et al.’s study of child sexual abuse material offenders [insoll2024factors], investigating the prevalence of help-seeking behavior (i.e., wishing to change) among deepfake creators to inform more effective interventions.
@@
Core members are also
Open Source and Academic Work. Core contributors to MrDeepFakes are technically skilled and willing excited to share their technical expertise on deepfake creationexpertise on creating deepfakes. Contrary to the beliefs of some open source developersof a SFW deepfake tool studied by Widder et al.  [widder2022ethicalai], denying technical support does not appear a viable mitigation for the abuse of deepfake creation toolsis not a viable intervention for deepfake tool abuse. Furthermore, significant sharing of academic work within the community shows that its members closely follow and utilize academic research, particularly from AI venues, highlighting the dual-use potential for many computing innovations. Similar to the OSS community, many instances of academic research have used technology neutrality and inevitability to justify the publication and dissemination of innovative developments. We acknowledge that the solution does not lie in silencing research findings or technological innovation, but we argue that researchers and developers should consider principles of “target hardening” in their own work to: (1) carefully choose how to distribute their workand , (2) rigorously evaluate and improve upon existing AI alignment methods, and, when possible, (23) preemptively design and integrate safeguards to promote the intended use and discourage the abuse of their researchtools.
@@ 7.2 Deepfake Abuse Mitigation is Adversarial
Technical solutions to for preventing deepfake creation are will likely be limited, as evidenced by adversarial the successful efforts of MrDeepFakes members to circumvent existing efforts by cloud GPU anti-abuse mechanisms added by cloud providers. While identifying platform-centered choke points (e.g., payment processors, hosting providers, ad providers) can reduce harm, as the security community has concluded for numerous abuse other ecosystems [franklin2007inquiry, han2022infra, krebs2010body], the endless emergence of alternative “bulletproof” service providers and cryptocurrencies has limited the efficacy of policy-centered mitigation (i.e., deplatforming)mitigations like deplatforming. In the face of complex security problems with no tractible, perfect solutions, we acknowledge emphasize the impact and importance of target hardening [redmiles2024friction]; however, we identify also argue that legislation may prove a promising approach to more effectively stem deepfake abuse through legislationfor curbing the creation of non-consensual sexual deepfakes.
@@
Banning Deepfake NCII.A MrDeepFakes staff member shared that the enactment of legislation internationally has already begun to make complicate monetizing sexual deepfakesmore difficult, implying that laws have significant sway over the deepfake economy (Section 5.4). In February 2024, following a push for federal legislation protecting against deepfake abuse [sarnoff2024taylor], forum members asked staff for their reflections on the implications of federally outlawing deepfake abuse, to which a MrDeepFakes staff member replied:
@@
If deepfakes are banned, I could finally do something else with my life without feeling bad about quitting. I’ll also have time for creative hobbies that I can publicly share instead of making porn
@@
This suggests that laws are powerful enough to disrupt behaviors of Laws may dissuade even highly skilled, core members of the community. We argue that augmenting our existing mitigation approach by incorporating legislative protections has significant potential to more effectively potential to help stem deepfake abuse.
Setting Ethical Norms for Deepfakes.We also observed that MrDeepFakes ’s rules defining celebrities as acceptable targets of deepfake NCII were laxly enforced, indicating a difference between the community’s written rules and normative behaviors in practice. Echoing Timmerman et al’̇s findings [timmerman2023studying], we presented evidence that members’ beliefs regarding the ethics of deepfake NCII creation ranged widely (Section 5.2). Such dissonance even within the MrDeepFakes community illustrates the broader state of flux surrounding the ethics of deepfake content. This liminality presents an opportunity for policy impact beyond raising deepfake creation cost disincentives, through guiding education and legal process approaches. We encourage policymakers to update sexual education curriculum to confront technology-mediated sexual interactions; through early education regarding concepts such as sexual safety, consent, and bodily autonomy in digital-physical ecosystems, we can begin to shape resilient norms surrounding the creation and consumption of deepfake content that proactively mitigate its harms.
Seeking Justice & Reducing Harm.Finally, we preemptively broach legal considerations in light of the mainstreaming of deepfake content and increasing commoditization of deepfake creation tools, as evidenced by the growth of MrDeepFakes and its discussion of limited-data tools. We reiterate Pfefferkorn’s recommendations regarding justice pathways for victims and perpetrators of deepfake NCII: that legislators consider the implications of criminal prosecution for preventing deepfake IBSA—especially regarding children. Due to rising popularity of commodity deepfake NCII tools among teens [maiberg24survey], blanket criminal charges for perpetrators also affect children; Pfefferkorn argues that criminal prosecution of children is counterproductive, unjustifiable, and causes lasting harm [pfefferkorn2024teens]. Therefore, we argue that legislators carefully balance the trade-offs of retributive and restorative justice pathways and acknowledge the nuances of criminalization when children are also known perpetrators.
8 Conclusion
In this work, we analyzed the MrDeepFakes platform since its inception and characterized the community’s motivations and deepfake creation process. Our investigation surfaces significant growth of deepfake NCII creation and consumption. We further find that beyond nonconsensual sexual gratification and power as understood motivations for IBSA perpetrators, our qualitative analysis reveals novel motivations for buyers of deepfake NCII: the perception of deepfakes as art and the distribution of new deepfakes as a community contribution. We document the requirements for deepfake creation, the challenges it poses, and the respective community-driven solutions. Our work sheds light on the rapidly growing, complex ecosystem around deepfake NCII, and we highlight opportunities to design proactive and reactive interventions to better mitigate deepfake IBSA.
EthicsEthical Considerations
@@ 8 Conclusion
In this paper, we analyze MrDeepFakes, a publicly available platform for hosting, buying, and sellingpublic platform for buying, selling, and hosting sexual deepfakes. Our work is in the spirit of significant prior work by the security community [motoyama11analysis, tseng20tools, wei2024understanding, soska15measuring, gamage2022deepfakes]. However, our context is significantly differentfrom prior workdifferent, and we carefully evaluated the ethical implications of our specific research. We considered many stakeholders when evaluating the ethics of our work: the the potential harm to several stakeholders: the depicted and discussed targets, the MrDeepFakes platformitself, its users, depicted and discussed targets, and the researchers who studied the platform given its potentially harmful contentand the researchers themselves. We discuss the ethical implications of our research choices for each stakeholder below: .
The Platform.
Targets of Sexual Deepfakes. To avoid drawing undue attention to any targets mentioned in the forum, we redact any personally identifying information in the quotes we report and do not name any targeted individuals in the text.
@@
The MrDeepFakes Platform. MrDeepFakes’ terms of service prohibit using site content for commercial endeavors, redistributing content, and sending users unsolicited emailsmessages, but they do not prohibit scraping or automated data collectionotherwise prohibit scraping. The site’s robots.txt only disallows crawling by the Internet Archive (User-agent: ia_archiver). All of the information we scraped from the website scraped information is accessible to anyone with an account.   [^4] Account creation is free and does not require e-mail verification; site operators have also stated that users are not obligated to provide “real” information. (Account creation is free and does not require e-mail verification; site operators have also stated that users are not obligated to provide “real” information.) We acknowledge that in scraping the website, we may have increased server load and bandwidth on the part of the platform; however, we minimized the number of requests made to the platform, and our request volume is relatively low compared to the organic activity of tens of thousands of accounts that the platform services. We stress that we , such that our study has minimal impact on the platform.
We did not collect or store any video or image data, only collecting relevant metadata required to answer our research questions.
@@
We did consider We considered whether to publish the name of the platform itself in this paper, as similar prior work opted to not reveal the name of the platforms in question platform names as to not draw potential attention to its their existence [tseng20tools]. However, we decided to name the MrDeepFakes platform as it has been named directly in prior work [timmerman2023studying, popova2020reading] and is the top search result for celebrity sexual deepfakes. We hope that naming this platform can enable future research and potential regulation on the issue of further research and policy on sexual deepfake creation and mainstreaming.
Users of the platform.
In examining user data, we collected only publicly available data, and
@@
Users of MrDeepFakes . In examining public user profiles, the only potentially personally identifying information we collected was the account username and the site-assigned user ID. Accounts are pseudonymous: there are no fields for “real” names. We never attempted to deanonymize any userin our dataset , and we did not interact with any community members in any fashion community members (e.g., via direct messages or public posts). In our report, we take careful note to not include any users’care to exclude any user’s account details when reporting quotes and opt to only report report only aggregate statistics to protect user privacy.
@@
To avoid drawing undue attention to any targets mentioned in the forum, we redact and paraphase any personally identifying information in the quotes we reported and elect not to name any target in the textPer a publication requirement from the USENIX Security program committee, we paraphrased quotes to mitigate the possible re-identification of quoted users.
@@
Researchers. Given that this paper required qualitative analysis of text-based forum data, the researchers involved in these tasks were directly exposed to potentially disturbing content that might cause significant disturbing text discussion and deepfake content that could cause psychological or emotional harms. distress. The research team was aware of these potential harms but recognized that exposure to this content was unavoidable to gain domain knowledge, understand the affordances of rendered webpages, and verify collected data. We ultimately made an informed decision to move forward with our study because we believed in the timeliness and importance of investigating this ecosystem. To mitigate these issuesharms, the researchers had regular check-ins regarding their psychological well being, well-being, worked together as a team to process MrDeepFakes content, had access to therapy and counseling services, and took breaks as needed.
@@
We are committed to sharing our data and code with other researchers at academic, non-commercial institutions seeking to conduct future work. However, given the sensitive nature of the data and disturbing nature of the contentdata, we will not make this data available publicly. not make this data available publicly. We note that the MrDeepFakes site content is available on the open Internet and that motivated actors can easily access the contentfor themselves. However, we do not want to further enable malicious actors seeking to use MrDeepFakes data to potentially harm others. We are committed to sharing our data and our codebooks with the Artifact Evaluation committee to ensure our artifacts meet the USENIX Open Science standards.
Appendix A MrDeepFakes Forum mapSection Map
The hierarchical map of the forum’s sections.